# A Domain-specific Language and Architecture for Detecting Process Activities from Sensor Streams in IoT

### v1

@@ 3 Related Work
Two significant works discuss challenges and opportunities of using IoT data together with BPM systems, highlighting the issue of event abstraction from low-level data to BPM-related data [3, 19], and more specific challenges associated with using data from sensors in this context [5]. Mangler et al. conclude that the relevance of sensors and dependencies among sensors contributing to the detection of BPM activities are difficult to be derived automatically and often need to involve domain experts. Moreover, sensor data associated with the execution of activities of the same type might be subject to variations due to external factors and parameters. Especially data from continuous sensors is affected by variations, which requires discretizations, pre-processing and abstraction steps often relying on domain expertise [5].
3.1 Activity Detection from Sensor Data
@@
In [20], the authors discuss the monitoring of business processes by automatically generating complex event processing (CEP) queries associated with the lifecycle transitions of control flow elements. Complementary to this work, [21] proposes to automatically learn and generate CEP rules for business activity detection from historical event traces. While our intent of monitoring business processes is similar, the authors rely on data from a BPM system to be available to trigger and log corresponding events. In our work, we first need to perform an event abstraction step to go from the IoT data to BPM-related execution data, which may then serve as a basis for subsequent process mining steps. Note that the actual analysis of these process data using process mining techniques is out of scope of our work. Various approaches investigate detecting and deriving BPM-related information from low-level sensor data.
@@
Janssen et al. discuss the discovery of process models from sensor event data in [9], based on machine learning (ML) with average quality results. Automatically derived, recurring patterns in sensor data enriched with context information are used to discover human activities and habits in [10, 22]. In [11] the authors derive the operations of a smart car for process mining based on a statistical analysis of the car’s sensor data. The authors in [23] focus on the detection of activities in manual processes based on different modalities. A plethora of works propose trained, supervised models to detect human activities from specific sensors (e.g., wearables [24, 25] or cameras [26]). We aim to develop a more general approach that is applicable to any kind of activity–manual or automated–and arbitrary sensor data. A generic approach towards the activity detection from IoT data has been proposed in [27], which uses one prototypical execution of an activity associated with its sensor data to generate a detection service. While being able to derive abstracted, process-related information, all of the aforementioned approaches are not able to handle variations in the underlying sensor data well as they are overfitting, and most of them do not consider existing domain expertise in their models. Moreover, they are only used for post-mortem detection and classification of activities. With our work, we aim to also enable online process analysis–streaming process mining [12]–as feedback at runtime is crucial for activity executions and decision making in IoT [28]. Furthermore, even though the given problem might be suitable to be approached by ML [29], the training of corresponding models is often expensive and leads to large, monolithic and inexplicable models, which are hard to maintain and adjust. As we aim to derive abstracted, high-level events from CPS sensors to inform about (business) process activity executions, we will focus on pattern-based sensor data processing following a more light-weight pipes and filters approach. The representative IoT systems serving as running examples in this work (cf. Sections 2.1 and 2.2) indicate that relevant patterns to detect starts and and ends of process activities might be referring to events from a rather small number of sensors and their combinations [30], which would make a pattern matching approach feasible, especially for runtime detections [31]. We will also discuss how more complex sensors (e.g., cameras [32]) and sensor networks can be integrated into our activity detection approach following a pre-processing and abstraction phase to represent more realistic IoT settings [33].
@@ 3.2 Domain-specific Languages for Activity Detection
We acknowledge that there exists a large corpus of work related to DSLs [34, 35] and their engineering [36, 37], also in the context of CPS [38, 39, 40] and IoT [41, 42]. DSL-based approaches supporting the general modeling of components, their interactions, and data flow in CPS and IoT systems are presented in [43, 40, 41], and with a special focus on monitoring of IoT components in [42]. More specialized DSLs in the these contexts feature for example the modeling of hazards and risks in CPS [44] and job scheduling in CPS [39]. While these languages feature aspects that we can adapt to model the IoT system and its components (e.g., the grouping of sensors and actuators), none of the DSLs is suitable to perform the necessary event abstraction step and they do not support BPM-related concepts, which is also the case for many approaches that aim at reverse engineering of low-level logs (e.g., containing interaction traces) to derive higher level activities (e.g., in software development processes [45]). Complex event processing (CEP) has proven to be a suitable technology for abstracting low-level data to higher level events, also in the BPM domain, according to Soffer et al. [46]. CEP-based platforms usually feature their own SQL-like language to specify the event processing rules and applications. The authors in [47] argue tat CEP languages are still too complex and require technical expertise to be usable by domain experts. Various approaches propose simplifications of these language via abstraction and graphical composition. In [48] Boubeta-Puig et al. propose a model-driven approach to facilitate the user-friendly design and composition of generic CEP patterns using a visual representation. A visual framework for data flow programming in IoT based on CEP is introduced in [49].
@@
Even though it is a relevant research problem [3], none of the aforementioned DSLs and approaches share the goal of detecting generic business process activities from arbitrary sensors. The approaches closest to this goal can be found in [50] proposing a DSL for human activity detections in smart homes and in [51] with a visual language to detect human activities based on IoT data. While we adapt some patterns and rules from these language, we acknowledge that there is a need to develop a DSL which is agnostic to a specific application domain and works with arbitrary sensor data. Moreover, targeting domain experts with a rather technical background, we will focus on developing textual DSL of efficient sensor pattern specification, rather than a visual notation [52].
@@ 4 Domain-specific Language: Radiant
We present the textual domain-specific language (DSL) Radiant, which enables domain experts to specify event abstraction patterns in sensor data to detect activity executions. Radiant has been developed following the principles and concepts for domain-specific event processing languages laid out in [53] and in [54]. In general, we followed the design science research methodology [55] in developing the DSL as main artifact. We closely aligned the requirements and concepts to be integrated into the meta-model and concrete syntax of the language with experts from the domains of smart manufacturing and smart healthcare in multiple iterations of meetings and workshops. Here we discussed process activities to be detected, sensors to be used, and generic IoT sensor patterns that may indicate process events and activity executions. From these discussions, we developed the meta-model of Radiant that combines concepts from the BPM and IoT domains for activity detection as depicted in Figure 5 and we integrate the identified generic patterns into its concrete syntax (cf. Section 4.2). Radiant is designed to be a domain-specific language for detecting process- level activities from sensor data in arbitrary IoT systems and IoT domains that feature sensors as main data sources.
4.1 Meta-model
@@
Figure 5: Meta-model of Radiant (green: concepts supported by the DSL; blue: IoT configurations in external YAML file; white: additional BPM concepts)
@@
The meta-model of the Radiant language is based on the event-centric meta-model for IoT driven process monitoring presented in  [56]. As depicted in Figure 5, a Process contains one or more Activities, which we aim to detect. For simplicity reasons, we do not consider nested process structures (i.e., subprocesses) in the current version of the meta-model. A corresponding extension is feasible to be implemented. An Activity is associated with exactly one startPattern, exactly one endPattern, and an arbitrary number of intermediatePatterns that are used for detecting the activity. The start and end patterns delimit the occurrence of an activity by a Start Event and End Event as Process Events, which are to be emitted denoting their occurrence (cf. Figures 1 and 3) and used as basis for process mining in subsequent data analyses. A Pattern is specified by one or more Conditions that have to be fulfilled for it to be detected. A Condition refers to the value(s) of one Sensor. These concepts are integrated in the syntax of Radiant and will be explained in more detail in Section 4.2.
@@
Furthermore, the meta-model supports the representation of the IoT system, which is necessary for technical configurations of the runtime system (cf. Section 4.6) and DSL support (e.g., auto-completion and sensor value verification). An IoT System is composed of one or more Stations, which aggregate their associated Sensors. A Station is also associated with a Source representing the event stream that contains the station’s sensor values as event payload. To avoid overloading the domain expert with these technical configuration issues, we decided to separate the configuration of the IoT system from the concepts integrated into concrete syntax of the DSL. The configurations have to be provided by an IoT engineer in the form of an externally composed YAML file as explained in Section 4.3 and are automatically loaded when using the DSL in an IDE.
4.2 Syntax
@@
Based on the presented meta-model, we now explain the concrete syntax of Radiant in more detail. Note that the following listings presenting the grammar of Radiant are based on the grammar language of Langium [^1] https://langium.org/, which was used to implement the DSL.
@@
The main concepts of Radiant that were adapted from the BPM domain are that of a Process and an Activity [13]. The process represents the context to which an activity belongs. An arbitrary number of activities (at least one) to be detected can be contained in a process (cf. Listing LABEL:lst:syntax-process), their ordering is of no relevance and there are no dependencies among activities. At this point it is important to be able to associate a detected activity with the corresponding process. Conformance checking (i.e., verifying the correct sequence of activity executions w.r.t. to a process model [6]) is subject to later analysis steps in process mining.
@@
Patterns
An Activity is characterized by a number of Patterns within the IoT sensor data that have to be detected, most importantly a startPattern and an endPattern, which are both mandatory (cf. Listing LABEL:lst:syntax-activity) as they represent the relevant process-level events to be used for process mining (cf. Section 2). An arbitrary number of intermediatePatterns can be specified in between the start and end pattern to make the activity detection more precise and robust. This might be needed to resolve potential ambiguities when more than one activity is characterized by the same patterns [57]. If the domain expert is aware of relevant sensor (change) patterns in the data, they should be specified as intermediate pattern to a reasonable extend (i.e., no overspecification by including dozens or hundreds of patterns) to make the detection robust and progress trackable. Intentional underspecification is also a valid approach to be followed by the domain expert to cope with potential variations in the underlying sensor data [5]. The patterns defined to be part of an activity are all dependent in the order of their specification from the start pattern to the end pattern, i.e., the activity detection assumes that the occurrence of the start pattern indicates its start, then the intermediate patterns from first to last, and then the end pattern to indicate its end.
1Activity:
@@ 2 Activity name=ID ’:’
Each pattern consists of one or more Conditions specifying when the pattern is detected (cf. Listing LABEL:lst:syntax-pattern). As shown in more detail in Listing LABEL:lst:syntax-patterncase we support conjunctions and disjunctions of an arbitrary number of conditions. Hereby, each Case keyword indicates that the conditions should be considered to be linked by a logical OR. Otherwise, and within one Case statement, the list of conditions specified for a pattern are considered to be linked by a logical AND.
@@ 14 Case:’
The most important concepts adapted from IoT are sensors indicating the values to be analyzed and their location in a specific IoT system or device (e.g., production station) [2], which is relevant for locating the specific event stream to be processed in our setup. These concepts are the base of a condition, which is displayed in Listing LABEL:lst:syntax-base. This base is complemented by one mandatory condition type from the ones listed in Table 1. Radiant currently supports these 10 different types as they cover the detection of all process activities we encountered in our use cases and discussions with domain experts. This list can be extended easily. Conditions that indicate a specific change (types ChangeCondition, ChangingCondition, IncreasingCondition, DecreasingCondition) can optionally be extended with a time constraint (cf. Listing LABEL:lst:syntax-time) defining a time-based window within which the change must happen to fire a higher-level event [58, 59].
@@
2 In station=ID sensor sensor=ID ConditionType
@@ 4.3 IoT Configurations
When working with rather low-level sensor data from IoT (e.g., as shown in Listings LABEL:lst:sorting and LABEL:lst:donor), the requirement of sensor discretization emerged to make the specification of the conditions more user friendly and more robust against variations in continuous data (e.g., by considering ranges in the sensor discretization, as opposed to exact values that need to be met) [5]. Moreover, available sensors, their data types, value ranges, and associations with IoT devices have to be pre-configured to make the DSL more usable and support the domain expert with auto-suggestions, validations (e.g., sensor values being out of range) and code completions–increasing productivity [60]. While these aspects might also be part of the DSL, we decided to treat them as a separate, more technical concern and externalize it into an external YAML-based configuration file following the structure and attributes of entities presented in the meta-model (cf. Section 4.1). Listing LABEL:lst:config shows parts of this sensor configuration for the sorting machine station. Here we see the definition of a template for the motor speeds (lines 1–8) with state abstractions and min/max values that can be reused in the specification of the actual motor sensors (line 24) as all motors in the smart factory have this same behavior. The listing also contains an exemplary discretization of the Integer values emitted from the color sensor (lines 18–21).
@@
The sources and sinks of the sensor event streams need to be configured, as well. Our goal is to detect activity executions at runtime from streams of IoT sensor data. We assume that typical messaging systems and brokers (e.g., MQTT, RabbitMQ, Apache Kafka, etc.) are available as sources to emit these sensor events and we can consume them for analysis in a publish-subscribe manner [61]. The configurations to connect to these systems and map the event data to the internal event processing data models are also part of the external configuration file containing the sensor configurations. For the sinks of the event processing, we suggest to emit the process-level start and end events from the activity detection in a standardized format for subsequent process analysis (e.g., in XES format [62]) on one stream per process instance [63]. Listing LABEL:lst:config also presents an exemplary source stream configuration (lines 31–43) to access the event streams from the sorting machine station of the smart factory via MQTT, which is referred to via the ID in the source parameter of the sensor configurations (cf. listing LABEL:lst:config, line 13). In addition, the listing contains the schema definition of the JSON-based messages received on this event stream (lines 38–43), which is used to map the payload to the internal event model.

### v2

@@ 3 Related Work
Two significant works discuss challenges and opportunities of using IoT data together with BPM systems, highlighting the issue of event abstraction from low-level data to BPM-related data [janiesch2020internet, brzychczy2025process], and more specific challenges associated with using data from sensors in this context [mangler2024internet]. Mangler et al. conclude that the relevance of sensors and dependencies among sensors contributing to the detection of BPM activities are difficult to be derived automatically and often need to involve domain experts. Moreover, sensor data associated with the execution of activities of the same type might be subject to variations due to external factors and parameters. Especially data from continuous sensors is affected by variations, which requires discretizations, pre-processing and abstraction steps often relying on domain expertise [mangler2024internet]. These observations confirm our focus on enabling the domain expert to define and configure the IoT-based event abstraction (cf. RQ1).
3.1 Event Abstraction from Sensor Data
@@
In [DBLP:conf/icsoc/BackmannBH0W13], the authors discuss the monitoring of business processes by automatically generating complex event processing (CEP) queries associated with the lifecycle transitions of control flow elements. Complementary to this work, [mousheimish2016autocep] proposes to automatically learn and generate CEP rules for business activity detection from historical event traces. While our intent of monitoring the execution of business processes is similar, the authors rely on data from a BPM system to be available to trigger and log corresponding events. Assuming that BPM systems are not always available in IoT systems, we first need to perform an event abstraction step to go from the IoT data to BPM-related execution data [diba2020extraction], which is the main functional objective of our work.
Various approaches investigate detecting and deriving BPM-related information from low-level sensor data. Janssen et al. discuss the discovery of process models from sensor event data in [janssen2020process], based on machine learning (ML). Automatically derived, recurring patterns in sensor data enriched with context information are used to discover human activities and habits in [di2022vamos, di2023cvamos]. In [beyel2024analyzing] the authors derive the operations of a smart car for process mining based on a statistical analysis of the car’s sensor data. The authors in [rebmann2019enabling] focus on the detection of activities in manual processes based on different modalities. Large language models (LLMs) are investigated in the IoTMiner framework by [brzychczy2025iot]. Following some pre-processing steps, the LLMs are prompted to classify segmented data from IoT based on pre-defined activity types. While these works fulfill the main functional objective of event abstraction and are applicable to generic IoT data (Objective 4), they only partially include domain expertise (e.g., in the LLM prompt), and they introduce the need for a training data being already available, potentially with additional labels (e.g., when using ML). None of them has demonstrated online capabilities as they rely on a post-mortem offline analysis for classification (Objective 3).
@@
Even though the given problem might be suitable to be approached by ML [wang2019deep], the training of corresponding models requires existing data and labels, is often resource-intensive and leads to large, monolithic and inexplicable models, which are hard to maintain and adjust, especially by domain experts who might not be familiar with ML (Objective 1). Many works propose trained, supervised models to detect human activities from sensors (e.g., wearables [GarciaCejaBCG14, cornacchia2016survey], smartphones [yin2015human] or cameras [aggarwal2014human]). These works are only applicable to specific types of sensors, they only work in offline settings, and they require extensive existing training data and ML models that do not foster the understandability of the event abstraction (Objective 1) [nweke2019data]. Even ML-models for sensor-based activity detection that are tailored and heavily optimized for edge computing scenarios (TinyML) are 130 kb in size or much larger [zhou2025efficient, sharma2024efficient] and additionally require a model-serving infrastructure. We aim to develop a more general, light-weight approach that is applicable to any kind of activity–manual or automated–and arbitrary IoT sensor data (Objectives 3 and 4). Most of the ML-based approaches do not consider existing domain expertise in their models, they need to rely on data and might incorporate some expertise in the provided labels. Our goal is to treat the expertise of the domain experts regarding sensors and sensor patterns as first class citizens, without the need to have a corpus of training data already available. Moreover, many related ML-based approach are only capable of post-mortem detection and classification of activities. With our work, we also enable online process analysis–streaming process mining [burattin2022streaming]–as feedback at runtime is crucial for activity executions and decision making in IoT (Objective 2). As we aim to derive abstracted, high-level events from CPS sensors to inform about (business) process activity executions, we will focus on pattern-based sensor data processing following a more light-weight pipes and filters style [khezemi2024systematic] (Objective 3). The IoT systems serving as running examples in this work (cf. Sections 2.3 and 2.4) indicate that relevant patterns to detect starts and ends of process activities might be referring to events from a rather small number of sensors and their combinations [kirikkayis2023integrating], which would make a pattern matching approach based on domain knowledge feasible, especially for runtime detections [klein2013using]. We will also discuss how more complex sensors (e.g., cameras [malburg2021object]) and sensor networks can be integrated into our activity detection approach following a pre-processing and abstraction phase to represent more realistic IoT settings [seiger5165943online].
@@ 3.2 Domain-specific Languages for Activity Detection
We acknowledge that a large corpus of work related to DSLs [kosar2016domain, iung_systematic_2020] and their engineering [krahn2010monticore, kleppe2008software], also in the context of CPS [butting2023towards, wood2021triton, pradhan2015chariot] and IoT [salihbegovic2015design, erazo2022domain], already exists. The main intent of a DSL is to allow domain experts to specify some kind of data or control-related operations based on the available domain knowledge, not requiring extensive programming skills (Objective 1) or existing data and resources for model training (Objective 3) [hudak1997domain].
@@
DSL-based approaches supporting the general modeling of components, their interactions, and data flow in CPS and IoT systems are presented in [tichy2020experiences, pradhan2015chariot, salihbegovic2015design], and with a special focus on monitoring of IoT components in [erazo2022domain]. More specialized DSLs in the these contexts feature for example the modeling of hazards and risks in CPS [petzold_pasta_2023], job scheduling in CPS [wood2021triton], or simplified data flow and rule modeling for specific IoT-based condition and location monitoring use cases [choosumrongdevelopment, phasinam2025real]. These languages feature aspects that we can adapt to model the IoT system and its components (e.g., the grouping of sensors and actuators) in a sensor-agnostic way (Objective 4) and they foster understandability as patterns and rules become explicit (Objective 1). However, none of the DSLs is able to fulfill the functional requirement to perform the necessary event abstraction step and they do not support BPM-related concepts, which is also the case for many approaches that aim at abstracting low-level logs (e.g., containing interaction traces) to derive higher level activities (e.g., in software development processes [caldeira2016software]). Complex event processing (CEP) has proven to be a suitable, light-weight technology (Objective 3) for abstracting low-level data to higher level events with runtime capabilities (Objective 2), also in the BPM domain according to Soffer et al. [soffer2019event]. CEP-based platforms usually feature their own SQL-like language to specify the event processing rules and applications. The authors in [rosa-bilbao_model-driven_2022] argue that CEP languages are still too complex and require technical expertise to be usable by domain experts. Various approaches propose simplifications of these language via abstraction and graphical composition. In [boubeta-puig_model-driven_2014] Boubeta-Puig et al. propose a model-driven approach to facilitate the user-friendly design and composition of generic CEP patterns using a visual representation. A visual framework for data flow programming in IoT based on CEP is introduced in [gokalp2019visual].
@@
Even though it is a relevant research problem [janiesch2020internet], none of the aforementioned DSLs and approaches share the functional objective of detecting generic business process activities from arbitrary sensors. The approaches closest to this goal can be found in [negrete_ramirez_dsl-based_2021] proposing a DSL for human activity detections in smart homes and in [barricelli_visual_2017] with a visual language to detect human activities based on IoT data. While we adapt some patterns and rules from these languages, we acknowledge that there is a need to develop a DSL which is agnostic to a specific application domain, which works with arbitrary sensor data and independent of specific activity types, and which features BPM-related concepts as first class citizens to abstract the data for subsequent process mining analysis. Moreover, targeting domain experts with a rather technical IoT background, we will focus on developing textual DSL of efficient sensor pattern specification, rather than a visual notation [gronninger2014textbased].
@@ 4 Domain-specific Language: Radiant
Addressing the research question RQ1, we present the textual domain-specific language (DSL) Radiant, which allows domain experts to specify event abstraction patterns in sensor data to detect activity executions as the main functional objective. Radiant has been developed following the principles and concepts for domain-specific event processing languages laid out in [bruns2014ds] and in [karsai2014design].
4.1 Meta-model of Radiant
@@
Figure 6: Meta-model of Radiant in UML (green: concepts supported by the DSL; blue: IoT configurations in external YAML file; white: additional BPM concepts)
@@
The meta-model of the Radiant language is based on the event-centric meta-model for IoT driven process monitoring presented in [franceschetti2023event]. As depicted in Figure 6, a Process contains one or more Activities [dumas2013fundamentals], which we aim to detect. For simplicity reasons, we do not consider nested process structures (i.e., subprocesses) in the current version of the meta-model. An Activity is associated with exactly one startPattern, exactly one endPattern, and an arbitrary number of intermediatePatterns. The start and end patterns delimit the occurrence of an activity by a Start Event and End Event as Process Events, which are to be emitted denoting their occurrence (cf. Figures 2 and 4) and used as basis for process mining in subsequent data analyses. A Pattern is specified by one or more Conditions that have to be fulfilled for it to be detected. A Condition refers to the value(s) of one Sensor.
@@
Furthermore, the meta-model supports the representation of the IoT system, which is necessary for technical configurations of the runtime system (cf. Section 4.6) and DSL support. An IoT System is composed of one or more Stations, which aggregate their associated Sensors. A Station is also associated with a Source representing the data stream that contains the station’s sensor values as payload. To avoid overloading the domain expert with these technical configuration issues, we decided to separate the configuration of the IoT system from the concepts integrated into concrete syntax of the DSL.
4.2 Syntax of Radiant
@@
Based on the presented meta-model, we now explain the concrete syntax of Radiant in more detail. Note that the following listings presenting the grammar of Radiant are based on the grammar language of Langium [^3] https://langium.org/, which was used as a modern, extensible and flexible framework to implement the DSL [bork2023language].
@@
The main concepts of Radiant that were adapted from the BPM domain are that of a Process and an Activity. In Radiant, the process represents the context to which an activity belongs. An arbitrary number of activities (at least one) to be detected can be contained in a process (cf. Listing LABEL:lst:syntax-process). As the process concept here only contextualizes an activity, the activity ordering is of no relevance and there are no dependencies among activities. Process discovery and conformance checking are subject to later analysis steps in process mining, which are out of scope of this work.
@@
3
4Activity:
@@ 5 Activity name=ID ’:’
Patterns
An Activity is characterized by a number of Patterns within the IoT sensor data that have to be detected, most importantly a startPattern and an endPattern, which are both mandatory (cf. Listing LABEL:lst:syntax-process) as they represent the relevant process-level events to be used for process mining (cf. Section 2). An arbitrary number of intermediatePatterns can be specified in between the start and end pattern to make the activity detection more precise and robust. This might be needed to resolve potential ambiguities when more than one activity is characterized by the same patterns [franceschetti2023characterisation]. If the domain expert is aware of relevant sensor patterns in the data, they should be specified as intermediate pattern to make the detection robust and progress trackable. The patterns defined to be part of an activity are all dependent in the order of their specification from the start pattern, to intermediate patterns, to the end pattern–allowing for stateful processing.
@@
Each pattern consists of one or more Conditions specifying when the pattern is detected based on the different types (Start, Intermediate, or End). As shown in detail in Listing LABEL:lst:syntax-patterncase, we support conjunctions and disjunctions of an arbitrary number of conditions. Hereby, each Case keyword indicates that the conditions should be considered to be linked by a logical OR. Otherwise, and within one Case statement, the list of conditions specified for a pattern are considered to be linked by a logical AND.
@@ 17 Case:’
The most important concepts adapted from IoT are sensors indicating the values to be analyzed and their location in a specific IoT system or device (e.g., production station) [bauer2013iot], which is relevant for locating the specific data stream to be processed in our setup. These concepts are the base of a condition, which is displayed in Listing LABEL:lst:syntax-base. This base is complemented by one mandatory condition type from the ones listed in Table 1. Radiant currently supports these 10 different types as they cover the detection of all process activities we encountered in our use cases and discussions with domain experts. This list can be extended easily. Conditions that indicate a specific change (types ChangeCondition, ChangingCondition, IncreasingCondition, DecreasingCondition) can optionally be extended with a time constraint defining a time-based window within which the change must happen to fire a higher-level event [jain2008towards, etzion2010event]. The ChangingCondition refers to any change in the specific sensor value, which can be made more coarse-grained via discretization (cf. Section 4.3).
@@
2 In station=ID sensor sensor=ID ConditionType (time_constraint+=TimeConstraint)*
3
4TimeConstraint:
5 within amount=INT time_unit=ID;
@@ 4.3 IoT Configurations
When working with rather low-level sensor data from IoT (e.g., as shown in Listings LABEL:lst:sorting and LABEL:lst:donor), the requirement of sensor discretization emerged to make the specification of the conditions more user friendly and more robust against variations in continuous data (e.g., by considering ranges in the sensor discretization, as opposed to exact values that need to be met) [mangler2024internet]. Using descretization, change-related conditions from Table 1 can also be made more coarse-grained to account for sensors that show a high frequency of insignificant changes (e.g., environmental sensors with high precision); thus, to indicate changes in the patterns of the discretized sensor values that are more relevant for activity detection.
Moreover, available sensors, their data types, value ranges, and associations with IoT devices have to be pre-configured to make the DSL more usable and support the domain expert with auto-suggestions, validations (e.g., sensor values being out of range) and code completions–increasing productivity [kosar2008preliminary]. While these aspects might also be part of the DSL, we decided to treat them as a separate, more technical concern and externalize it into an external YAML-based configuration file following the structure and attributes of entities presented in the meta-model (cf. Section 4.1). Listing LABEL:lst:config shows parts of this sensor configuration for the sorting machine station. Here we see the definition of a template for the motor speeds (lines 1–8) with state abstractions and min/max values that can be reused in the specification of the actual motor sensors (line 24) as all motors in the smart factory have this same behavior. The listing also contains an exemplary discretization of the Integer values emitted from the color sensor (lines 18–21).
@@
The sources and sinks of the sensor data streams need to be configured, as well. Our goal is to detect activity executions at runtime from streams of IoT sensor data. We assume that typical messaging systems and brokers (e.g., MQTT, RabbitMQ, Apache Kafka, etc.) are available as sources to emit these sensor events and we can consume them for analysis in a publish-subscribe manner [corral2020stream]. The configurations to connect to these systems and map the event data to the internal event processing data models are also part of the external configuration file. For the sinks of the event processing, we suggest to emit the process-level start and end events from the activity detection in a standardized format for subsequent process analysis (e.g., in XES format [gunther2014xes]) on one stream per process instance [burattin2020mqtt]. Listing LABEL:lst:config also presents an exemplary source stream configuration (lines 31–43) to access the event streams from the sorting machine station of the smart factory via MQTT, which is referred to via the ID in the source parameter of the sensor configurations (cf. listing LABEL:lst:config, line 13). In addition, the listing contains the schema definition of the JSON-based messages received on this event stream (lines 38–43), which is used to map the payload to the internal event model.